u(2,0) = 6 5 Solve the equation by using implicit FDM(,)-4 for u u(3t) 2 At-1
Answer #1

lul ou i-u ox At ot substitute these values in equation(1) Lu Ax Ax substitute ΔΧ_Δ1-1substitute i = Lj = 0 in equation (2) u 2u +1+2(0- 2.0 substitute i-2, j-0 In equation (2) 2,1 2.0 2-2(6)+6+2(2-6)=1121-6 → 2